OpenStudy (anonymous):

will medal also fan anyone who helps i have 2 questions

OpenStudy (nathalyn):

ok

OpenStudy (anonymous):

what is \[3^{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

cube root of nine square root of nine cube root of 3 square root of 3

OpenStudy (anonymous):

@NathalyN help

OpenStudy (anonymous):

@amoodarya @Bird_113 @Cubi-Cal

OpenStudy (nathalyn):

its not 1 or 2

OpenStudy (anonymous):

k by the way \[\frac{ 2 }{ 3 }\] is an exponent

OpenStudy (nathalyn):

√3 = ± 1.7320508

OpenStudy (nathalyn):

for ther 1 one

OpenStudy (anonymous):

ok

OpenStudy (nathalyn):

2^3 = 8

OpenStudy (anonymous):

does anyone else think its a

OpenStudy (anonymous):

@Mehek14 help

OpenStudy (anonymous):

@KakeFace

OpenStudy (anonymous):

@Keigh2015 help

OpenStudy (anonymous):

@Mehek14 @Keigh2015 @KakeFace

Mehek (mehek14):

this is how you do it Example: \(z^\dfrac{5}{3} = 3\sqrt{z^5}\)

OpenStudy (anonymous):

oh

Mehek (mehek14):

so your denominator is 3 it goes outside the radical \[^3\sqrt{3^2}\]

Mehek (mehek14):

and \(3^2=9\)

Mehek (mehek14):

\[^3\sqrt{9}\]

OpenStudy (keigh2015):

Are they answered?

OpenStudy (anonymous):

ok so \[\sqrt[2]{3^{3}}\]

OpenStudy (anonymous):

i have one more

Mehek (mehek14):

no the other way \[^3\sqrt{3^2}\]

OpenStudy (anonymous):

correct steps to find the value of \[8^{\frac{ 1 }{ 3 }}\]?

OpenStudy (anonymous):

ok

Mehek (mehek14):

\[\sqrt[3]{8^1}\]

Mehek (mehek14):

\(8^1=8\) \[\sqrt[3]{8}\]

Mehek (mehek14):

now find the cube root of 8

OpenStudy (anonymous):

ok

Mehek (mehek14):

what do you get?

OpenStudy (anonymous):

2

Mehek (mehek14):

that's how you solve it \(\ddot\smile\)

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw

OpenStudy (anonymous):

fan me back?

OpenStudy (anonymous):

@Mehek14 fan me back :)?

sammixboo (sammixboo):